# Refusal Direction is Universal Across Safety-Aligned Languages

### v1

@@ Abstract
Refusal mechanisms in large language models (LLMs) are essential for ensuring safety. Recent research has revealed that refusal behavior can be mediated by a single direction in activation space, enabling targeted interventions to bypass refusals. While this is primarily demonstrated in an English-centric context, appropriate refusal behavior is important for any language, but poorly understood. In this paper, we investigate the refusal behavior in LLMs across 14 languages using PolyRefuse, a multilingual safety dataset created by translating malicious and benign English prompts into these languages. We uncover the surprising cross-lingual universality of the refusal direction: a vector extracted from English can bypass refusals in other languages with near-perfect effectiveness, without any additional fine-tuning. Even more remarkably, refusal directions derived from any safety-aligned language transfer seamlessly to others. We attribute this transferability to the parallelism of refusal vectors across languages in the embedding space and identify the underlying mechanism behind cross-lingual jailbreaks. These findings provide actionable insights for building more robust multilingual safety defenses and pave the way for a deeper mechanistic understanding of cross-lingual vulnerabilities in LLMs. [^1] We make our code publicly available at https://github.com/mainlp/Multilingual-Refusal. Equal Contribution.
@@ 1 Introduction
LLMs are increasingly deployed across a wide range of real-world applications (Kaddour et al., 2023; Yang et al., 2024; Raza et al., 2025). To ensure their safe use, LLMs are expected to exhibit a refusal mechanism, the ability to obey to non-harmful request but refuse harmful, unethical, or policy-violating requests (Bai et al., 2022). This capability is typically instilled via reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Christiano et al., 2017; Dai et al., 2024) and other alignment strategies (Yuan et al., 2024; Wallace et al., 2024; Xu et al., 2024).
Despite these efforts, LLMs remain vulnerable to jailbreak attacks, including adversarial prompt engineering (Wei et al., 2023; Zou et al., 2023; Liu et al., 2024a; Tao et al., 2024), where carefully crafted inputs trigger unsafe outputs, and targeted fine-tuning (Yang et al., 2023; Lermen et al., 2024; Zhan et al., 2024), which undermines safety constraints through parameter updates. Notably, cross-lingual jailbreaks have emerged as a growing concern (Yong et al., 2024; Li et al., 2024; Deng et al., 2024), where prompts in non-English languages bypass refusal mechanisms that succeed in English, raising critical questions about the multilingual refusal mechanism in LLMs.
@@ 2 Related Work
In AI safety research, various efforts have been made to prevent LLMs from responding to malicious queries. Notable approaches include supervised fine-tuning (SFT) (Bianchi et al., 2024) and reinforcement learning from human feedback (RLHF) (Bai et al., 2022). To evaluate the effectiveness of these safety measures, researchers have developed comprehensive safety evaluation datasets. While these datasets initially focused on English (Zou et al., 2023; Mazeika et al., 2024; Xie et al., 2025), recent work has expanded to include multilingual evaluations, revealing concerning vulnerabilities in non-English contexts (Shen et al., 2024; Yong et al., 2024; Wang et al., 2024). Furthermore, researchers have begun investigating the internal mechanisms that enable LLMs to recognize and refuse harmful requests. Studies examining model representations have identified specific “refusal directions” in the embedding space (Arditi et al., 2024; Marshall et al., 2025). However, these mechanistic interpretability studies have predominantly focused on English, leaving cross-lingual aspects of refusal mechanisms largely unexplored. This paper addresses this gap by investigating how refusal mechanisms function across different languages.
@@ 3.1 Refusal Direction Extraction
The candidate refusal vectors are obtained by collecting the difference-in-means vectors across all layers at final instruction token positions, such as the [/INST] token for Llama2 (Touvron et al., 2023). The most effective refusal vector is then identified by evaluating the reduction in refusal behavior after ablating each candidate from the residual stream and choosing the one with the most reduction in refusal behavior, measured by the drop of refusal score after ablating the vector (Arditi et al., 2024). The refusal score calculates the probability difference between refusal-associated tokens $\mathcal{R}$ (e.g., ‘Sorry’, ‘I’ for English) and non-refusal tokens $\𝒱\mathcal{V}\backslash\mathcal{R}$, calculated at the initial token position of the model’s generation:
Refusal $\textit{Refusal Score}=\log\left(\sum_{t\in\mathcal{R}}p_{t}\right)-\log\left(\sum_{t\in\mathcal{V}\backslash\mathcal{R}}p_{t}\right)$ (3)
@@ 3.2 Removing or Adding Refusal Behavior
$\mathbf{x}^{\prime}_{l}\leftarrow\mathbf{x}_{l}-\mathbf{\hat{r}}_{l}\mathbf{\hat{r}}_{l}^{\top}\mathbf{x}_{l}$ (4)
@@ 4 Not All Languages are Safety-Aligned
English-centric safety alignment has been shown to generalize poorly to other languages (Yong et al., 2024; Li et al., 2024; Deng et al., 2024). To assess cross-lingual jailbreak vulnerability, we show compliance rates across 14 languages using three instruction-tuned models: Llama3.1-8B-Instruct, Qwen2.5-7B-Instruct, and gemma-2-9B-Instruct. Each model is tested on 572 translated harmful prompts per language (see dataset details in §5.1), with responses translated back into English and evaluated using WildGuard (Han et al., 2024). As shown in Table 1, all models exhibit varying levels of susceptibility across languages.
@@ 5.2 Cross-lingual Transfer of English Refusal Vectors
We evaluate a diverse set of models spanning multiple sizes, including Yi (Young et al., 2025), Qwen2.5 (Qwen et al., 2025), Llama-3 (Grattafiori et al., 2024), and Gemma-2 (Gemma Team et al., 2024), to ensure that our findings generalize across different model families and scales. We use Wildguard (Han et al., 2024) to classify whether the model refuses or complies with the queries. In the case of non-English queries and responses, we first translate them back into English before feeding them to Wildguard for classification. To investigate whether refusal vectors derived from English are transferable to other languages, we ablated these vectors from the residual stream as described in §3.2. We then measure the compliance rate to harmful queries before and after ablation. The results are presented in Figure 1.
Our first key finding is that English-derived refusal vectors lead to a substantial increase in harmful compliance across all evaluated models and safety-aligned languages. Even models that initially demonstrate strong multilingual safety, such as gemma-2-9B-it and Qwen2.5-14B-Instruct, can be successfully jailbroken post-ablation, highlighting that even the most robust multilingual safety mechanisms can be compromised by targeting a direction derived solely from English data.
@@ 6 Exploring the Geometry of Refusal in LLMs
Across all models and language pairs, we observe that the refusal directions – computed between harmful and harmless embeddings – are approximately parallel across languages. This alignment confirms the strong cross-lingual transferability discussed in §5 and suggests a shared internal axis of harmfulness. To quantify this observation, we compute the cosine similarity between each language’s refusal direction and the difference-in-means vectors of every other language across all post-instruction token positions [^6] The post-instruction token positions in Llama3.1-8B-Instruct are: “¡—eot_id—¿”, “¡—start_header_id—¿”, “assistant”, “¡—end_header_id—¿”, “ n n”, corresponding to positions -5 to -1, respectively. Empirically, positions -5 and -1 yield the most effective refusal extraction, consistent with our heatmap results. and decoder layers. The resulting heatmap, shown in Figure 4 for Llama3.1-8B-Instruct [^7] Heatmaps for other models are provided in §A.3.2., reveals consistently high cross-lingual similarity, at the token and layer position where the refusal vector was extracted for the source language. These results suggest that LLMs encode refusal signals in a structurally aligned and language-agnostic manner.
However, visualizations in Figure 3 also reveal a key vulnerability: while harmful and harmless samples form clearly separated clusters in English, the separation is substantially less distinct in non-English languages, especially in the Llama3.1-8B-Instruct model (cf. Figure 3(a), 3(b), 3(c), 3(d)). This weak clustering diminishes the strength of refusal signals, particularly in underrepresented or non-Latin-script languages. Jailbroken samples (red) always lie in the intermediate region between harmful and harmless clusters, indicating that the model struggles to decisively classify them.
@@
To quantify the clustering difference between English and non-English languages, we calculate the Silhouette Score (Rousseeuw, 1987) to assess clustering quality, evaluating how well harmful and harmless embeddings align with their respective clusters. A higher Silhouette Score means clearer separation of harmful and harmless content. As shown in Table 2, the Silhouette Scores in English are consistently higher across all models, confirming that harmful and harmless samples are more cleanly separated in English compared to other languages. These results quantitatively validate the insights from the PCA visualizations, highlighting the degradation of clustering quality in multilingual settings, although the universal refusal direction has been learned.
@@
Overall, while instruction-tuned models consistently learn a universal refusal direction, they fail to establish robust boundaries between harmful and harmless prompts in many languages. This insufficient separation is a key factor behind cross-lingual jailbreak vulnerabilities.
@@ A.2 Refusal Tokens
Table 4 presents the refusal tokens identified for each model across different languages. These tokens represent the most frequent sentence-initial tokens that appear distinctively when models refuse harmful requests compared to their responses to harmless prompts. The analysis reveals consistent patterns within language families—models typically begin refusals with first-person pronouns ("I", "{CJK}UTF8gbsn我", "Ich", "Я") or polite expressions ("{CJK}UTF8min申し訳" in Japanese, "{CJK}UTF8mj죄" in Korean).
@@
Qwen2.5-7B I, As {CJK}UTF8gbsn我 Es, Ich [Uncaptioned image] Mo {CJK}UTF8min申し, 私は Я, И {CJK}UTF8mj죄
gemma2-9B I, It {CJK}UTF8gbsn我, 很 Ich [Uncaptioned image] Mi, Emi {CJK}UTF8min申し訳, 私は Я, Из {CJK}UTF8mj죄
Llama-3.1-8B I {CJK}UTF8gbsn我 Ich [Uncaptioned image] I {CJK}UTF8min私は Я {CJK}UTF8mj저
@@ A.3.2 Refusal direction similarity heatmap
We present cross-lingual heatmaps for Qwen2.5-7B-Instruct and gemma-2-9B-it in Figures 6 and 7, respectively. Each heatmap visualizes the cosine similarity between the refusal direction extracted from a source language and the difference-in-means vectors of target languages across all decoder layers. Consistent with the observations from Llama3.1-8B-Instruct in Figure 4, both models demonstrate strong cross-lingual alignment of refusal signals, with similarity peaking around the extraction layers – approximately layers 15–29 in Qwen2.5-7B-Instruct and layers 19–23 in gemma-2-9B-it. These results further support the conclusion that multilingual refusal directions are structurally aligned and language-agnostic across models.
@@ A.4 Ablation Results for Other Languages
The ablation results demonstrate distinct patterns in cross-lingual generalization depending on the source language of refusal vectors. Figures 8(a)-8(d) show that refusal vectors extracted from safety-aligned languages (Japanese, Korean, Russian) exhibit strong generalization across multiple target languages. When ablating Japanese-derived vectors, compliance rates increase substantially across most languages for all three models, indicating that safety mechanisms learned in Japanese effectively transfer to other well-aligned languages. Similar patterns emerge for Korean and Russian vectors, suggesting that refusal representations in safety-aligned languages capture generalizable safety concepts. In contrast, Figure 8(d) reveals different behavior for Yoruba-derived vectors. The poor generalization observed across most languages reflects Yoruba’s limited safety alignment in the evaluated models, particularly evident in Qwen2.5-7B and Llama-3.1-8B, where baseline compliance rates are already low. Notably, Yoruba vectors show relatively better performance on Gemma-2-9B, consistent with this model’s superior safety alignment in Yoruba compared to the other models, as shown in Table 1. These findings indicate that cross-lingual transfer of safety mechanisms is contingent upon the source language’s alignment quality. Well-aligned languages produce refusal vectors that effectively generalize across the multilingual safety landscape, while poorly aligned languages yield vectors with limited transferability.

### v2

@@ Abstract
Refusal mechanisms in large language models (LLMs) are essential for ensuring safety. Recent research has revealed that refusal behavior can be mediated by a single direction in activation space, enabling targeted interventions to bypass refusals. While this is primarily demonstrated in an English-centric context, appropriate refusal behavior is important for any language, but poorly understood. In this paper, we investigate the refusal behavior in LLMs across 14 languages using PolyRefuse, a multilingual safety dataset created by translating malicious and benign English prompts into these languages. We uncover the surprising cross-lingual universality of the refusal direction: a vector extracted from English can bypass refusals in other languages with near-perfect effectiveness, without any additional fine-tuning. Even more remarkably, refusal directions derived from any safety-aligned language transfer seamlessly to others. We attribute this transferability to the parallelism of refusal vectors across languages in the embedding space and identify the underlying mechanism behind cross-lingual jailbreaks. These findings provide actionable insights for building more robust multilingual safety defenses and pave the way for a deeper mechanistic understanding of cross-lingual vulnerabilities in LLMs. [^1] We make our code publicly available at https://github.com/mainlp/Multilingual-Refusal.Equal Contribution.The work was done prior to joining Amazon.
@@ 1 Introduction
LLMs are increasingly deployed across a wide range of real-world applications (Kaddour et al., 2023; Yang et al., 2024; Raza et al., 2025). To ensure their safe use, LLMs are expected to exhibit a refusal mechanism, the ability to obey to non-harmful request but refuse harmful, unethical, or policy-violating requests (Bai et al., 2022). This capability is typically instilled via reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Christiano et al., 2017; Dai et al., 2024) and other alignment strategies (Yuan et al., 2024; Wallace et al., 2024; Xu et al., 2024a).
Despite these efforts, LLMs remain vulnerable to jailbreak attacks, including adversarial prompt engineering (Wei et al., 2023; Zou et al., 2023; Liu et al., 2024a; Tao et al., 2024), where carefully crafted inputs trigger unsafe outputs, and targeted fine-tuning (Yang et al., 2023; Lermen et al., 2024; Zhan et al., 2024), which undermines safety constraints through parameter updates. Notably, cross-lingual jailbreaks have emerged as a growing concern (Yong et al., 2024; Li et al., 2024a; Deng et al., 2024), where prompts in non-English languages bypass refusal mechanisms that succeed in English, raising critical questions about the multilingual refusal mechanism in LLMs.
@@ 2 Related Work
In AI safety research, various efforts have been made to prevent LLMs from responding to malicious queries. Notable approaches include supervised fine-tuning (SFT) (Bianchi et al., 2024) and reinforcement learning from human feedback (RLHF) (Bai et al., 2022). To evaluate the effectiveness of these safety measures, researchers have developed comprehensive safety evaluation datasets. While these datasets initially focused on English (Zou et al., 2023; Mazeika et al., 2024; Xie et al., 2025), recent work has expanded to include multilingual evaluations, revealing concerning vulnerabilities in non-English contexts (Shen et al., 2024; Yong et al., 2024; Wang et al., 2024). Furthermore, researchers have begun investigating the internal mechanisms that enable LLMs to recognize and refuse harmful requests. Studies examining model representations (Xu et al., 2024b; Li et al., 2024b) and have identified specific “refusal directions” in the embedding space (Arditi et al., 2024; Marshall et al., 2025). However, these mechanistic interpretability studies have predominantly focused on English, leaving cross-lingual aspects of refusal mechanisms largely unexplored. This paper addresses this gap by investigating how refusal mechanisms function across different languages.
@@ 3.1 Refusal Direction Extraction
The candidate refusal vectors are obtained by collecting the difference-in-means vectors across all layers at final instruction token positions, such as the [/INST] token for Llama2 (Touvron et al., 2023). The most effective refusal vector is then identified by evaluating the reduction in refusal behavior after ablating each candidate from the residual stream and choosing the one with the most reduction in refusal behavior, measured by the drop of refusal score after ablating the vector (Arditi et al., 2024). The refusal score calculates the probability difference between refusal-associated tokens $\mathcal{R}$ (e.g., ‘Sorry’, ‘I’ for English) and non-refusal tokens $\mathcal{V}\backslash\mathcal{R}$, calculated at the initial token position of the model’s generation:
$\textit{Refusal Score}=\log\left(\sum_{t\in\mathcal{R}}p_{t}\right)-\log\left(\sum_{t\in\mathcal{V}\backslash\mathcal{R}}p_{t}\right)$ (3)
@@ 3.2 Removing or Adding Refusal Behavior
$\mathbf{x}^{\prime}_{l}\leftarrow\mathbf{x}_{l}-\mathbf{\hat{r}}\mathbf{\hat{r}}^{\top}\mathbf{x}_{l}$ (4)
@@ 4 Not All Languages are Safety-Aligned
English-centric safety alignment has been shown to generalize poorly to other languages (Yong et al., 2024; Li et al., 2024a; Deng et al., 2024). To assess cross-lingual jailbreak vulnerability, we show compliance rates across 14 languages using three instruction-tuned models: Llama3.1-8B-Instruct, Qwen2.5-7B-Instruct, and gemma-2-9B-Instruct. Each model is tested on 572 translated harmful prompts per language (see dataset details in §5.1), with responses translated back into English and evaluated using WildGuard (Han et al., 2024). As shown in Table 1, all models exhibit varying levels of susceptibility across languages.
@@ 5.2 Cross-lingual Transfer of English Refusal Vectors
We evaluate a diverse set of models spanning multiple sizes, including Yi (Young et al., 2025), Qwen2.5 (Qwen et al., 2025), Llama-3 (Grattafiori et al., 2024), and Gemma-2 (Gemma Team et al., 2024), to ensure that our findings generalize across different model families and scales. We use Wildguard (Han et al., 2024) to classify whether the model refuses or complies with the queries. In the case of non-English queries and responses, we first translate them back into English before feeding them to Wildguard for classification.
To investigate whether refusal vectors derived from English are transferable to other languages, we ablated these vectors from the residual stream as described in §3.2. We then measure the compliance rate to harmful queries before and after ablation. The results are presented in Figure 1. Our first key finding is that English-derived refusal vectors lead to a substantial increase in harmful compliance across all evaluated models and safety-aligned languages. Even models that initially demonstrate strong multilingual safety, such as gemma-2-9B-it and Qwen2.5-14B-Instruct, can be successfully jailbroken post-ablation, highlighting that even the most robust multilingual safety mechanisms can be compromised by targeting a direction derived solely from English data.
@@ 5.3 Refusal Vectors from Non-English Languages
5.4 Reducing Compliance Rate by Adding Refusal Vectors
Figure 3: Compliance rates to harmful queries before and after adding refusal vectors derived from EN, DE and TH. Vector addition decreases compliance across languages, demonstrating controllable refusal behavior.
To validate the controllability of refusal vectors, we performed the inverse manipulation by adding refusal vectors to the residual stream, as described in §3.2. We extracted refusal vectors from three languages: English, German, and Thai. Figure 3 shows that vector addition consistently decreases compliance rates across languages, demonstrating the inverse effect of ablation.
All three source languages achieve strong reduction in compliance across most languages and models, with the majority reaching near-zero compliance rates after addition. For instance, adding English vectors reduces compliance to 0% for most languages in gemma-2-9b-it, with Yoruba dropping from 58% to 28%. German and Thai vectors demonstrate comparable effectiveness, with most languages achieving 0% compliance across the three models.
Notably, Yoruba exhibits different behavior compared to other languages: it maintains higher residual compliance rates across all three source languages, though the pattern varies by model. This aligns with our previous observation in the ablation experiments: refusal direction is only transferable across safety-aligned languages. It is hard to induce or remove refusal in a language where the concept of safety alignment does not exist.
The consistent effectiveness of refusal vectors across diverse source languages combined with the symmetry between ablation and addition, demonstrates that refusal vectors enable controllable modulation of safety behavior across languages. This finding supports our universality hypothesis: refusal mechanisms learned from any safety-aligned language can transfer effectively across the multilingual space.
@@ 6 Exploring the Geometry of Refusal in LLMs
Figure 4: PCA visualizations of multilingual harmful and harmless representations in the refusal extraction layer. Top: Llama3.1-8B-Instruct. Middle: Qwen2.5-7B-Instruct. Bottom: gemma-2-9B-it. Arrows indicating refusal directions per language.
@@
Across all models and language pairs, we observe that the refusal directions – computed between harmful and harmless embeddings – are approximately parallel across languages. [^6] Arrows are illustrative and not strictly drawn from centroids. To quantify this observation, we compute the cosine similarity between each language’s refusal direction and the difference-in-means vectors of every other language across all post-instruction token positions [^7] The post-instruction token positions in Llama3.1-8B-Instruct are: “¡—eot_id—¿”, “¡—start_header_id—¿”, “assistant”, “¡—end_header_id—¿”, “ n n”, corresponding to positions -5 to -1, respectively. Empirically, positions -5 and -1 yield the most effective refusal extraction, consistent with our heatmap results. and decoder layers. The resulting heatmap, shown in Figure 5 for Llama3.1-8B-Instruct [^8] Heatmaps for other models are provided in §A.3.2., reveals consistently high cross-lingual similarity, at the token and layer position where the refusal vector was extracted for the source language. These results suggest that LLMs encode refusal signals in a structurally aligned and language-agnostic manner.
However, visualizations in Figure 4 also reveal a key vulnerability: while harmful and harmless samples form clearly separated clusters in English, the separation is substantially less distinct in non-English languages, especially in the Llama3.1-8B-Instruct model (cf. Figure 4(a), 4(b), 4(c), 4(d)). This weak clustering diminishes the strength of refusal signals, particularly in underrepresented or non-Latin-script languages. Jailbroken samples (red) always lie in the intermediate region between harmful and harmless clusters, indicating that the model struggles to decisively classify them.
@@
To quantify the clustering difference between English and non-English languages, we calculate the Silhouette Score (Rousseeuw, 1987) to assess clustering quality, evaluating how well harmful and harmless embeddings align with their respective clusters. The Silhouette Score assesses clustering quality by combining two factors: (1) cluster compactness—how closely related a point is to others within the same cluster; and (2) between-cluster separation—how well the clusters are distinguished from each other. In our setting, we measure how well the representations of harmful and harmless queries form distinct clusters. As shown in Table 2, the Silhouette Scores in English are consistently higher across all models, confirming that harmful and harmless samples are more cleanly separated in English compared to other languages. These results quantitatively validate the insights from the PCA visualizations, highlighting the degradation of clustering quality in multilingual settings. For yo (Yoruba), the compact but poorly-separated harmful representations yield a moderate Silhouette Score despite the model’s safety misalignment, occasionally exceeding scores of better-aligned languages like de (German).
@@
Overall, while instruction-tuned models consistently learn a universal refusal direction, they fail to establish robust boundaries between harmful and harmless prompts in many languages at the refusal extraction layer. This insufficient separation is a key factor behind cross-lingual jailbreak vulnerabilities. It is important to note that our analysis centers around the refusal extraction layer, where the refusal behaviour is triggered. Recent work (Zhao et al., 2025) has shown that ’harmfulness’ and ’refusal’ are encoded differently in the model. Since the model needs to first identify the harmfulness concept before triggering the refusal behavior, our refusal vector analysis serves as a ‘downstream’ analysis of the ‘upstream’ harmfulness identification event.
@@ A.2 Refusal Tokens
Table 4 presents the refusal tokens identified for each model across different languages. These tokens represent the most frequent sentence-initial tokens that appear distinctively when models refuse harmful requests compared to their responses to harmless prompts. The analysis reveals consistent patterns within language families—models typically begin refusals with first-person pronouns ("I", "我", "Ich", "Я") or polite expressions ("申し訳" in Japanese, "죄" in Korean).
@@
Qwen2.5-7B I, As Es, Ich [Uncaptioned image] Mo 申し, 私は Я, И
gemma2-9B I, It 我, 很 Ich [Uncaptioned image] Mi, Emi 申し訳, 私は Я, Из
Llama-3.1-8B I Ich [Uncaptioned image] I 私は Я
@@ A.3.2 Refusal direction similarity heatmap
We present cross-lingual heatmaps for Qwen2.5-7B-Instruct and gemma-2-9B-it in Figures 7 and 8, respectively. Each heatmap visualizes the cosine similarity between the refusal direction extracted from a source language and the difference-in-means vectors of target languages across all decoder layers. Consistent with the observations from Llama3.1-8B-Instruct in Figure 5, both models demonstrate strong cross-lingual alignment of refusal signals, with similarity peaking around the extraction layers – approximately layers 15–29 in Qwen2.5-7B-Instruct and layers 19–23 in gemma-2-9B-it. These results further support the conclusion that multilingual refusal directions are structurally aligned and language-agnostic across models.
@@ A.4 Ablation Results for Other Languages
The ablation results demonstrate distinct patterns in cross-lingual generalization depending on the source language of refusal vectors. Figures 9(a)-9(d) show that refusal vectors extracted from safety-aligned languages (Japanese, Korean, Russian) exhibit strong generalization across multiple target languages. When ablating Japanese-derived vectors, compliance rates increase substantially across most languages for all three models, indicating that safety mechanisms learned in Japanese effectively transfer to other well-aligned languages. Similar patterns emerge for Korean and Russian vectors, suggesting that refusal representations in safety-aligned languages capture generalizable safety concepts. In contrast, Figure 9(d) reveals different behavior for Yoruba-derived vectors. The poor generalization observed across most languages reflects Yoruba’s limited safety alignment in the evaluated models, particularly evident in Qwen2.5-7B and Llama-3.1-8B, where baseline compliance rates are already low. Notably, Yoruba vectors show relatively better performance on Gemma-2-9B, consistent with this model’s superior safety alignment in Yoruba compared to the other models, as shown in Table 1. These findings indicate that cross-lingual transfer of safety mechanisms is contingent upon the source language’s alignment quality. Well-aligned languages produce refusal vectors that effectively generalize across the multilingual safety landscape, while poorly aligned languages yield vectors with limited transferability.
@@
A.5 Ablation Across Layers and Token Positions
Following Arditi et al. [2024], we conducted a comprehensive sweep across all layers and token positions across models and languages. We measured two key metrics: (1) KL divergence between original and ablated first token probability distributions, which quantifies the distributional shift caused by ablation, and (2) refusal score, which directly measures the model’s propensity to refuse harmful requests.
Figures 1015 present the results of this sweep analysis across gemma-2-9b-it, Llama-3.1-8B-Instruct, and Qwen2.5-7B-Instruct. Each figure shows the KL or Refusal Score change when ablating the candidate refusal vector extracted in the corresponding languages. The yellow highlighted regions indicate our final selected ablation targets. Due to KL filtering, the selected refusal vector exhibits relatively low KL divergence while achieving substantial reductions in refusal scores. This maximizes the attack’s effectiveness (high refusal score reduction) while minimizing unwanted side effects on the model’s general behavior (low KL divergence).
Table 8 shows the effect of KL filtering, where ablating selected refusal vector keep the model’s capability, measured by MMLU [Hendrycks et al., 2021], PPL on Wikitext [Merity et al., 2016], and ARC-C [Clark et al., 2018] while reducing the refusal rate dramatically.
Table 8: English Refusal Vector ablation result for Llama-3.1-8B-Instruct. Ablating the refusal vector removes refusal while maintaining the model capacity. Ablating other candidate diff-in-means vectors hurts the model capacity and fails to remove refusal completely. This is achieved by KL filtering and refusal score ranking.
Condition Refusal Rate MMLU PPL ARC-C
Before ablation 0.99 68.5 8.65 52.4
Ablate Refusal Vector 0.02 68.0 8.71 52.5
Ablate Random vector 1 (layer 12, token position 3) 1.0 65.8 9.17 49.6
Ablate Random vector 2 (layer 3, token position 3) 0.55 66.0 9.21 32.8
Figure 10: KL divergence between original and ablated first token probability distributions across layers and token positions for gemma-2-9b-it. Higher KL divergence indicates larger distributional changes from ablation, while lower values indicate minimal impact on model behavior. Yellow highlighted regions indicate the selected layers and tokens for refusal vector extraction.
Figure 11: Refusal score across layers and token positions for gemma-2-9b-it. Yellow highlighted regions correspond to the final selected ablation targets.
Figure 12: KL divergence between original and ablated first token probability distributions across layers and token positions for Llama-3.1-8B-Instruct. Yellow highlighted regions indicate the selected layers and tokens for refusal vector extraction.
Figure 13: Refusal score across layers and token positions for Llama-3.1-8B-Instruct. Yellow highlighted regions correspond to the final selected ablation targets.
Figure 14: KL divergence between original and ablated first token probability distributions across layers and token positions for Qwen2.5-7B-Instruct. Yellow highlighted regions indicate the selected layers and tokens for refusal vector extraction.
Figure 15: Refusal score across layers and token positions for Qwen2.5-7B-Instruct. Yellow highlighted regions correspond to the final selected ablation targets.